# Improving Temporal Understanding Logic Consistency in Video-Language Models via Attention Enhancement

### v1

@@ Abstract
Large language models (LLMs) often generate self-contradictory outputs, which severely impacts their reliability and hinders their adoption in practical applications. In video-language models (Video-LLMs), this phenomenon recently draws the attention of researchers. Specifically, these models fail to provide logically consistent responses to rephrased questions based on their grounding outputs. However, the underlying causes of this phenomenon remain underexplored. In this work, we adopt an interpretability-driven approach to analyze, statistically summarize, and intervention the potential factors of the phenomenon. We find that one of the primary reasons for the inconsistency in responses lies in the inability of cross-modal attention heads to effectively distinguish video tokens across different timestamps. To address this, we propose an attention enhancement method called Temporally Conditioned Attention Sharpening (TCAS), which constructs an enhancement objective based on attention distinctions to enhance the model’s temporal resolution capability, thereby improving its temporal understanding logic consistency. Experimental results demonstrate that our method significantly enhances the temporal logic consistency of Video-LLMs. Further interpretability analyses reveal that our method indeed improves the temporal discriminability of attention heads, validating our conclusions. Additionally, our method achieves performance improvements in general video temporal grounding tasks, highlighting that temporal logic consistency is a bottleneck in temporal understanding. By enhancing consistency, our method drives significant progress in video temporal understanding.
Improving Temporal Understanding Logic Consistency in Video-Language Models via Attention Enhancement
Chengzhi Li   Heyan Huang   Ping Jianthanks:   Corresponding author   Zhen Yang   Yaning Tian School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China {lichengzhi,hhy63,pjian,bityangzhen,1120222204}@bit.edu.cn
@@ 1 Introduction
In recent years, Multimodal Large Language Models (MLLMs) have achieved significant strides in visual-language understanding Yue et al. (2024); Ying et al. (2024). Among them, Video Large Language Models (Video-LLMs) excel by combining video-based visual information with the pre-trained knowledge of large language models, enabling better comprehension and generation of natural language related to video content Tang et al. (2025). These models have demonstrated outstanding performance in tasks like video question answering and captioning, driving progress in multimodal intelligence Fu et al. (2024). However, the redundancy in video data challenges large language models in fine-grained temporal understanding. To address this, many studies aim to enhance the temporal understanding Guo et al. (2025); Ren et al. (2024), particularly for fine-grained tasks like Video Temporal Grounding (VTG).
These works often rely on additional specialized temporal modules to enhance Video-LLMs by providing additional temporal information streams Guo et al. (2025); Ren et al. (2024); Hu et al. (2024). Jung et al. Jung et al. (2024) proposed a benchmark to assess the consistency of state-of-the-art (SOTA) Video-LLMs, revealing their significant limitations in fine-grained temporal understanding. Specifically, Jung et al. Jung et al. (2024) find that all Video-LLMs struggle to provide logically consistent responses to rephrased questions, highlighting a lack of true temporal comprehension in both the models and the enhancement methods.
Although many modular methods have been proposed to enhance temporal understanding, the underlying reasons for large language models’ limitations in this area remain underexplored. Jung’s findings Jung et al. (2024) further reinforce the necessity for deeper analysis. Our work begins with temporal inconsistency, constructing an interpretability-driven analysis to uncover its intrinsic causes and deeply exploring temporal understanding. To our knowledge, this is the first study to analyze the temporal understanding consistency of Video-LLMs from an interpretability perspective. We focus on the following research questions:
@@
2). How can we enhance the temporal understanding consistency of Video-LLMs based on the identified internal factors? In Section 3, we introduce Temporally Conditioned Attention Sharpening (TCAS), an optimization objective based on attention distributions. This compels the cross-modal attention heads to explicitly judge the relevance of temporally distributed information, improving the model’s temporal logic consistency. In Section 4, we validate the effectiveness across various experimental settings, demonstrating its ability to enhance temporal discriminability. Moreover, our method improves performance in general VTG tasks, highlighting temporal logic consistency as a key bottleneck.
@@ 2 Exploring Temporal Logic Consistency
We conduct experiments using the original grounding (G.), rephrased grounding (R-G.), and shifted grounding (S-G.) subsets of Charades-CON Jung et al. (2024). For each sample, we define the model’s Rephrase Consistency Score $c^{v}_{rg.}$ and Shift Consistency Score $c^{v}_{sg.}$ as $I^{v}_{ori}\times I^{v}_{rg.}$ and $I^{v}_{ori}\times I^{v}_{sg.}$, respectively, where $I^{v}_{ori}$, $I^{v}_{rg.}$, and $I^{v}_{sg.}$ represent the Intersection over Union (IoU) between the model predictions and the ground-truth timestamps for original, rephrased, and shifted grounding tasks.
This section focuses on analyzing the TimeChat-7b model Ren et al. (2024), which is a multimodal large language model used for long Video understanding and accurate time localization tasks, fusing temporal features and visual information through Sliding Video Q-Former and Timestation-aware frame encoder. In addition, for Generalizability of our conclusions, we analyze Qwen2.5-VL Bai et al. (2025) on the Event Order Judgment (EOJ) task in Section 2.4.
@@ 2.2 Statistical Analysis
$S^{h,v}_{disc.}=\frac{1}{|T_{e}|}\times\sum_{q\in T_{e}}\frac{\sum_{k\in V_{gt}}A^{h,v}_{q,k}}{\sum_{k\in V}A^{h,v}_{q,k}}$ (2)
$S^{v}_{disc.}=\frac{1}{|H_{t}|}\sum_{h\in H_{t}}S^{h,v}_{disc.}$ (3)
@@
For each video sample $v$, we calculate its consistency scores $c^{v}_{rg.}$, $c^{v}_{sg.}$ and attention discriminability score $S^{v}_{disc.}$. The Pearson correlation coefficient is computed by aggregating the values of these two metrics across all samples in the Charades-CON dataset. The discriminability scores distribution across different consistency sets is shown in Figure 4 as a violin plot, where we can find that $S_{disc.}$ and $c^{v}$ are positively correlated. The Pearson correlation coefficients are $0.4778$ and $0.4788$ for the rephrased grounding and the shifted grounding sets. This correlation coefficient is evidently statistically significant for Video-LLMs, which are inherently complex nonlinear systems. This result demonstrates a unambiguous positive correlation between the temporal discriminability of key heads and the model’s consistency performance. The scatter plot, the least-squares regression line, and more discussions are shown in Appendix A.2.
2.3 Causal Intervention
Statistical correlation alone is often insufficient to establish a causal relationship Pearl (2009). To further validate the causal relationship between attention discriminability and temporal logic consistency, a causal intervention Pearl (2009) study is conducted to determine whether improving attention discriminability leads to a rise in temporal logic consistency. Specifically, an intervention operation is designed that improves attention discriminability. By evaluating the model’s consistency performance after the intervention, we aim to confirm the causal relationship between attention discriminability and temporal logic consistency.
Specifically, the attention adjustment method is as follows: First, the multi-layer and multi-head attention distributions of the model during inference are obtained. Then, for the selected cross-modal attention heads with high attention scores, their query-to-key attention distributions are mixed with the ground-truth distribution within a specified intervention intensity $\alpha$. This process guides the model to focus more on the ground-truth temporal segments during inference. After the intervention, inference is performed again, and the changes in model outputs and attention distributions are compared to analyze the impact of the intervention on the model’s temporal understanding capabilities.
@@ 2.3 Causal Intervention
2.4 Generalizability Analysis
To validate the generalizability of our conclusions across other models and tasks, we extend the analysis to the Qwen2.5-VL Bai et al. (2025). Furthermore, we select the Event Order Judgment (EOJ) task as a representative one of complex temporal reasoning tasks Yu et al. (2023); Swetha et al. (2025), which involves temporal order judgment, allowing us to further investigate the model’s temporal capabilities in complex multi-event scenarios.
Figure 5: Violin plot of analysis results on EOJ task.
In the EOJ task, we define a new consistency score and an attention discriminability score to verify our findings. The distribution of attention discriminability scores for different ranges of consistency scores is shown as a violin plot in Figure 5. We observe a significant positive correlation (Pearson correlation coefficient is $0.447$), consistent with the trend found in TimeChat. This demonstrates that the relationship between attention discriminability and temporal logic consistency holds true for other leading Video-LLMs. The detailed implementation can be found in Appendix A.3.
@@ 3 Attention Enhancement
In the previous section, it is confirmed that the temporal discriminability of cross-modal attention heads is a key factor influencing the model’s temporal logic consistency. How can we enhance attention discriminability to improve temporal consistency? To achieve generalization and cross-architecture applicability, we formulate the attention discriminability enhancement as an optimization objective termed Temporally Conditioned Attention Sharpening Loss (TCAS Loss). The specific loss computation is detailed in Algorithm 1, where the $\textsc{SelectTopHeads}(H,S,t)$ means selecting the top $t$ attention heads from the set of all heads $H$ based on scores $S$.
@@
14:return $(\sum\limits_{h\in H_{t}}\sum_{q\in T^{h}_{valid}}\mathcal{L}^{h}_{q})/(\sum\limits_{h\in H_{t}}|T^{h}_{valid}|)$
We first identify the attention heads with the highest cross-modal scores and the text tokens that focus on specific video timestamps. This approach preserves the original contextual interaction capabilities of the attention heads. A contrastive learning loss is designed for the selected text query tokens to amplify the differences in attention scores across distinct timestamps, thereby compelling the attention heads to make explicit relevance judgments for temporally distributed information. We use weight $w_{ae}$ to balance the TCAS loss and next token prediction loss. Our method introduces no additional modules, ensuring compatibility with various transformer-based architectures Vaswani et al. (2023). Furthermore, it does not depend on specific text templates or video patterns, making it broadly applicable to diverse video-language tasks.
@@ 4.1 Experimental Setup
4.1.1 Models
We conduct experiments on the Qwen2.5-VL Bai et al. (2025), Video-LLaMA Zhang et al. (2023), and TimeChat Ren et al. (2024) models, which respectively represent Video-LLMs based on MLP projection, Video-LLMs based on Qformer, and time-aware models. More details about these models can be found in Appendix B.1.
4.1.2 Datasets
We conduct experiments on two widely-used VTG datasets: ActivityNet Captions Krishna et al. (2017) and Charades Gao et al. (2017). ActivityNet Captions comprises approximately 20k untrimmed YouTube videos with an average duration of 120 seconds, while Charades contains 9848 short videos with an average length of 30 seconds. Additionally, Charades-CON and ActivityNet-CON, introduced by Jung et al. Jung et al. (2024), serve as consistency benchmarks to ensure fair, transparent, and reliable comparisons.
4.1.3 Tuning Annotations
We select two different training Annotations. The first is a standard instruct annotation dataset for the VTG consistency task, called TimeIT, proposed by Ren et al. Ren et al. (2024). The second dataset is VTune, introduced by Jung et al. Jung et al. (2024), which has been shown to effectively enhance the consistency of Video-LLMs. Our method demonstrates improved results over baseline fine-tuning methods using both datasets, indicating that it does not rely on specific data sources. The dataset examples can be found in Appendix B.2.
4.1.4 Evaluation Metrics
Following previous works Jung et al. (2024); Ren et al. (2024), we evaluate grounding performance using IoU metrics, while measuring the consistency of the model responses through the accuracy rate. Details about the evaluation metrics can be found in Appendix B.3.
4.1.5 Implementation Details
We implement our method using PyTorch on a single NVIDIA A100 GPU. All settings are trained with the Adam optimizer (learning rate: $1\times 10^{-5}$, batch size: 4) for approximately 3 days. Qwen2.5-VL, TimeChat, and Video-LLaMA process 16, 96, and 8 frames respectively, all at $224\times 224$ spatial resolution.
@@ 4.1.5 Implementation Details
BM-DETR Jung et al. (2025) 59.4 38.3 49.6 30.6
Mr.BLIP Meinardus et al. (2024) 69.3 49.2 53.9 35.5
@@
HawkEye Wang et al. (2024) 58.3 28.8 34.7 17.7
VTG-LLM Guo et al. (2025) 57.2 33.4 - -
videollama (w. SFT) 37.1 20.1 34.3 19.1
videollama (w. TCAS) 37.5 20.1 35.2 20.0
Qwen2.5vl (w. SFT) 27.3 14.1 16.5 7.8
Qwen2.5vl (w. TCAS) 32.4 16.6 16.5 7.9
TimeChat Ren et al. (2024) 46.7 23.7 28.0 15.8
@@ 4.2.1 Baselines
We evaluate our method’s effectiveness by comparing it with several baseline models and approaches. For open-source models, direct comparison without fine-tuning would be unfair, as no existing modeling methods specifically target consistency enhancement. Therefore, we compare our approach with standard fine-tuning methods, specifically Supervised Fine-Tuning (SFT) Zhang et al. (2024), applied to several selected backbones. The results of other models without tuning include GPT-4o OpenAI (2024), Gemini-1.5 Flash Team (2024), and so on, listed in Appendix B.4.
Notably, some models, such as VTG-LLM Guo et al. (2025), focus on improving temporal understanding without addressing consistency issues. So we only compare their grounding performance, excluding consistency metrics.

### v2

# Understanding Temporal Logic Consistency in Video-Language Models through Cross-Modal Attention Discriminability

@@ Abstract
Large language models (LLMs) often generate self-contradictory outputs, which severely impacts their reliability and hinders their adoption in practical applications. In video-language models (Video-LLMs), this phenomenon recently draws the attention of researchers. Specifically, these models fail to provide logically consistent responses to rephrased questions based on their grounding outputs. However, the underlying causes of this phenomenon remain underexplored. In this work, we adopt an interpretability-driven approach to analyze, statistically summarize, and intervention the potential factors of the phenomenon. We find that one of the primary reasons for the inconsistency in responses lies in the inability of cross-modal attention heads to effectively distinguish video tokens across different timestamps. To address this, we propose an attention enhancement method called Temporally Conditioned Attention Sharpening (TCAS), which constructs an enhancement objective based on attention distinctions to enhance the model’s temporal resolution capability, thereby improving its temporal understanding logic consistency. Experimental results demonstrate that our method significantly enhances the temporal logic consistency of Video-LLMs. Further analyses reveal that our method indeed improves the temporal discriminability of attention heads, validating our conclusions. Additionally, our method even achieves performance improvements in general video temporal grounding tasks, suggesting that temporal logic consistency is an important factor in temporal understanding.
footnotetext: Corresponding author.
@@ 1 Introduction
In recent years, Multimodal Large Language Models (MLLMs) have achieved significant strides in visual-language understanding [27, 25]. Among them, Video Large Language Models (Video-LLMs) excel by combining video-based visual information with the pre-trained knowledge of large language models, enabling better comprehension and generation of natural language related to video content [20]. These models have demonstrated outstanding performance in tasks like video question answering and captioning, driving progress in multimodal intelligence [4]. However, the redundancy in video data challenges large language models in fine-grained temporal understanding. To address this, many studies aim to enhance the temporal understanding [6, 19], particularly for fine-grained tasks like Video Temporal Grounding (VTG).
These works often rely on additional specialized temporal modules to enhance Video-LLMs by providing additional temporal information streams [6, 19, 7]. Jung et al. [9] proposed a benchmark to assess the consistency of state-of-the-art (SOTA) Video-LLMs, revealing their significant limitations in fine-grained temporal understanding. Specifically, Jung et al. [9] find that all Video-LLMs struggle to provide logically consistent responses to rephrased questions, highlighting a lack of true temporal comprehension in both the models and the enhancement methods.
Although many modular methods have been proposed to enhance temporal understanding, the underlying reasons for large language models’ limitations in this area remain underexplored. Jung’s findings [9] further reinforce the necessity for deeper analysis. Our work begins with temporal inconsistency, constructing an interpretability-driven analysis to uncover its intrinsic causes and deeply exploring temporal understanding. To our knowledge, this is the first study to analyze the temporal understanding consistency of Video-LLMs from an interpretability perspective. We focus on the following research questions:
@@
2). How can we enhance the temporal understanding consistency of Video-LLMs based on the identified internal factors? In Section 4, we introduce Temporally Conditioned Attention Sharpening (TCAS), an optimization objective based on attention distributions. This compels the cross-modal attention heads to explicitly judge the relevance of temporally distributed information, improving the model’s temporal logic consistency. In Section 5, we validate the effectiveness across various experimental settings, demonstrating its ability to enhance temporal discriminability. Moreover, our method even improves performance in general VTG tasks, suggesting that temporal logic consistency is an important factor in temporal understanding.
@@
2 Related Work
2.1 Video-LLMs on VTG
Recent Video-LLMs excel in fine-grained temporal understanding tasks like VTG by independently encoding timestamps and interacting with video information to capture precise temporal segments, marking a significant improvement over traditional VTG methods by enabling diverse natural language outputs [6, 19, 7]. New benchmarks for fine-grained temporal understanding have been proposed to analyze their limitations [4, 2, 9]. For example, the consistency benchmark by Jung et al. [9] highlights response inconsistency. Unlike prior work, we focus on the internal representations of Video-LLMs rather than solely developing new benchmarks for evaluation, providing deeper insights into their temporal cognition.
2.2 Interpretability Analysis in MLLMs
In recent years, interpretability of LLMs has emerged as a key research focus to enhance model transparency and controllability [18]. However, the understanding of MLLMs remains limited [13]. Nikankin [15] showed that vision–language models rely on largely disjoint modality-specific circuits [3], with visual representations aligning to textual ones only in later layers. Venhoff [23] found that visual features gradually converge with language representations across layers, revealing an early-layer misalignment in adapter-based multimodal architectures. More recently, Yu and Ananiadou [26] conducted a mechanistic analysis of LLaVA and revealed that its visual reasoning process parallels the in-context learning dynamics of text-only LLMs.
The above work focuses on explaining how image concepts align with text representations. For temporal reasoning, Li et al. [12] probed visual features and identified the LLM decoder as the bottleneck, but did not explore the root causes of this limitation. From a consistency perspective, we offer a deeper explanation, attributing the decoder’s limitations to the insufficient discriminability of cross-modal attention heads.
@@ 3 Exploring Temporal Logic Consistency
We conduct experiments using the original grounding (G.), rephrased grounding (R-G.), and shifted grounding (S-G.) subsets of Charades-CON [9]. For each sample, we define the model’s Rephrase Consistency Score $c^{v}_{rg.}$ and Shift Consistency Score $c^{v}_{sg.}$ as $I^{v}_{ori}\times I^{v}_{rg.}$ and $I^{v}_{ori}\times I^{v}_{sg.}$, respectively, where $I^{v}_{ori}$, $I^{v}_{rg.}$, and $I^{v}_{sg.}$ represent the Intersection over Union (IoU) between the model predictions and the ground-truth timestamps for original, rephrased, and shifted grounding tasks.
This section focuses on analyzing the TimeChat-7b model [19], which is a multimodal large language model used for long Video understanding and accurate time localization tasks, fusing temporal features and visual information through Sliding Video Q-Former and Timestation-aware frame encoder. In addition, for Generalizability of our conclusions, we analyze Qwen2.5-VL [1] on the Event Order Judgment (EOJ) task in Appendix A.3.
@@ 3.2 Statistical Analysis
$S^{h,v}_{disc.}=\frac{1}{|T_{e}|}\times\sum_{q\in T_{e}}\frac{\sum_{k\in V_{gt}}A^{h,v}_{q,k}}{\sum_{k\in V}A^{h,v}_{q,k}}\text{,}$ (2)
$S^{v}_{disc.}=\frac{1}{|H_{t}|}\sum_{h\in H_{t}}S^{h,v}_{disc.}\text{,}$ (3)
@@
For each video sample $v$, we calculate its consistency scores $c^{v}_{rg.}$, $c^{v}_{sg.}$ and attention discriminability score $S^{v}_{disc.}$. The Pearson correlation coefficient is computed by aggregating the values of these two metrics across all samples in the Charades-CON dataset. The discriminability scores distribution across different consistency sets is shown in Figure 4 as a violin plot, where we can find that $S_{disc.}$ and $c^{v}$ are positively correlated. The Pearson correlation coefficients are $0.4778$ (with p-value $\mathbf{5.18\times 10^{-41}}\ll 0.05$) and $0.4788$ (with p-value $\mathbf{3.69\times 10^{-41}}\ll 0.05$) for the rephrased grounding and the shifted grounding sets. This correlation coefficient is evidently statistically significant for Video-LLMs, which are inherently complex nonlinear systems. This result demonstrates a unambiguous positive correlation between the temporal discriminability of key heads and the model’s consistency performance. The scatter plot, the least-squares regression line, and more discussions are shown in Appendix A.2.
@@
Figure 5: Attention patterns of key heads in the EOJ task.
Generalizability. To briefly explore the generalizability across other models and tasks, we extend our observation to the Qwen2.5-VL [1] on the Event Order Judgment (EOJ) task. As illustrated in Figure 5, we can observe that the text description tokens of events occurring at different times can generally establish appropriate focus on their corresponding temporal visual segments. This phenomenon implies the cross-model and cross-task generalizability of our findings. More detailed analysis is provided in Appendix A.3.
3.3 Intervention Verification
Statistical correlation alone is often insufficient to establish a causal relationship [17]. To further validate the causal relationship between attention discriminability and temporal logic consistency, a targeted intervention [17] study is conducted to determine whether improving attention discriminability leads to a rise in temporal logic consistency. Specifically, we intervene on attention for each query token $q\in T_{e}$ and head $h\in H_{t}$ via
$A_{q,V}=(1-\alpha)A_{q,V}^{orig}+\alpha A_{q,V}^{gt}.$ (4)
Here, $V$ is video tokens set, $A^{orig}$ is the original attention, and $A^{gt}$ is the ground-truth attention, a uniform distribution over tokens within the event timestamps. Targeted minor intervention isolates the effect of $S_{disc}$ excluding confounding factors. This process guides the model to focus more on the ground-truth temporal segments during inference.
By evaluating the model’s consistency performance after the intervention, we aim to confirm the causal relationship between attention discriminability and temporal logic consistency.
@@ 4 Attention Enhancement
In the previous section, it is confirmed that the temporal discriminability of cross-modal attention heads is a key factor influencing the model’s temporal logic consistency. How can we enhance attention discriminability to improve temporal consistency? To achieve generalization and cross-architecture applicability, we formulate the attention discriminability enhancement as an optimization objective termed Temporally Conditioned Attention Sharpening Loss (TCAS Loss).
To ensure cross-task generalization across various video-language tasks, Our TCAS Loss does not rely on ground-truth temporal labels to guide attention distribution optimization. This loss function employs a contrastive learning approach, compelling the model to make explicit relevance judgments for information at different timestamps, thereby enhancing the discriminability of attention heads in the temporal dimension.
@@
14:return $(\sum_{h\in H_{t}}\sum_{q\in T^{h}_{valid}}\mathcal{L}^{h}_{q})/(\sum_{h\in H_{t}}|T^{h}_{valid}|)$
@@
The specific loss computation is detailed in Algorithm 1, where the $\textsc{SelectTopHeads}(H,S,t)$ means selecting the top $t$ attention heads from the set of all heads $H$ based on scores $S$. We first identify the attention heads with the highest cross-modal scores and the text tokens that focus on specific video timestamps. This approach preserves the original contextual interaction capabilities of the attention heads. A contrastive learning loss is designed for the selected text query tokens to amplify the differences in attention scores across distinct timestamps, thereby compelling the attention heads to make explicit relevance judgments for temporally distributed information. We use weight $w_{ae}$ to balance the TCAS loss and next token prediction loss. Our method introduces no additional modules, ensuring compatibility with various transformer-based architectures [22]. Furthermore, it does not depend on specific text templates or video patterns, making it broadly applicable to diverse video-language tasks.
@@ 5.1 Experimental Setup
Models
We conduct experiments on the Qwen2.5-VL [1], Video-LLaMA [28], and TimeChat [19] models, which respectively represent Video-LLMs based on MLP projection, Video-LLMs based on Qformer, and time-aware models. More details about these models can be found in Appendix B.1.
Datasets
We conduct experiments on two widely-used VTG datasets: ActivityNet Captions [10] and Charades [5]. ActivityNet Captions comprises approximately 20k untrimmed YouTube videos with an average duration of 120 seconds, while Charades contains 9848 short videos with an average length of 30 seconds. Additionally, Charades-CON and ActivityNet-CON, introduced by Jung et al. [9], serve as consistency benchmarks to ensure fair, transparent, and reliable comparisons.
Tuning Annotations
We select two different training Annotations. The first is a standard instruct annotation dataset for the VTG consistency task, called TimeIT, proposed by Ren et al. [19]. The second dataset is VTune, introduced by Jung et al. [9], which has been shown to effectively enhance the consistency of Video-LLMs. Our method demonstrates improved results over baseline fine-tuning methods using both datasets, indicating that it does not rely on specific data sources. The dataset examples can be found in Appendix B.2.
Evaluation Metrics
Following previous works [9, 19], we evaluate grounding performance using IoU metrics, while measuring the consistency of the model responses through the accuracy rate. Details about the evaluation metrics can be found in Appendix B.3.
Implementation Details
We implement our method using PyTorch on a single NVIDIA A100 80GB GPU. All settings are trained with the Adam optimizer (learning rate: $1\times 10^{-5}$, batch size: 4) for approximately 3 days. Qwen2.5-VL, TimeChat, and Video-LLaMA process 16, 96, and 8 frames respectively, all at $224\times 224$ spatial resolution.
@@
BM-DETR [8] 59.4 38.3 49.6 30.6
Mr.BLIP [14] 69.3 49.2 53.9 35.5
@@
HawkEye [24] 58.3 28.8 34.7 17.7
VTG-LLM [6] 57.2 33.4 - -
TimeChat [19] 46.7 23.7 28.0 15.8
@@ 5.2.1 Baselines
We evaluate our method’s effectiveness by comparing it with several baseline models and approaches. For open-source models, direct comparison without fine-tuning would be unfair, as no existing modeling methods specifically target consistency enhancement. Therefore, we compare our approach with standard fine-tuning methods, specifically Supervised Fine-Tuning (SFT) [29], applied to several selected backbones. The results of other models or backbones are listed in Appendix B.4.